Can you help me with the following problem
Mr. López is responsible for the purchase of wine boxes for the Casa Blanca restaurant. Periodically choose a test box (12 bottles per box) to determine if the sealing process is adequate. For this test, randomly select four bottles from the box to taste the wine. If a box contains two bottles of wine in poor condition, calculate the probability that precisely one of them appears in Mr. López's sample.
Answer #1

P(precisely one of them appears in Mr. López's sample) =P(one from 2 in poor condition and 3 from 10 of good condition) =(2C1)*(10C3)/(12C4) =2*120/495=16/33 =0.4848